OpenStudy (goformit100):

If the cartesian coordinates of a point are given by (2, y) and its polar coordinates are (r, 30°), determine y and r. 4. Two points in a plane have polar coordinates (2.50 m, 30.0°) and (3.80 m, 120.0°). Determine (a) the cartesian coordinates of these points and (b) the distance between them.

Parth (parthkohli):

You know that the rectangular coordinates for polar coordinates \((r,\theta)\) are given by \((r\cos\theta, r\sin\theta)\).

Parth (parthkohli):

Use that fact, maybe?

OpenStudy (marco26):

x=rcos(theta) , y=rsin(theta)

Parth (parthkohli):

\[r\cos30^{\circ} = 2\]What is \(\cos(30^{\circ})\)?

OpenStudy (goformit100):

root 3 by 2

Parth (parthkohli):

Correct.\[r \times \dfrac{\sqrt{3}}{2} = 2\]Find \(r\).

OpenStudy (goformit100):

4 by root 3

Parth (parthkohli):

Nice! Can you find \(y\) now? It's \(r\sin(30^{\circ}) = \dfrac{4}{\sqrt3} \times 0.5\)

OpenStudy (goformit100):

Thankyou Sir

Parth (parthkohli):

Koi nahi re! X-D